Solar radiation enters the Earth's atmosphere with a portion being scattered by clouds and aerosols.

Processing, archiving and distributing Earth science data
at the NASA Langley Research Center

TES Instrument Operational Status

TES Instrument Operational Status

Thursday, September 21, 2017

UPDATE:  (09/21/2017)

On September 15, 2017, the Aura Tropospheric Emission Spectrometer (TES) Interferometer Control System (ICS) translator regained full motion.  Recovery activities were ongoing since the stall occurred on July 29, 2017.  TES resumed science collection on September 17, 2017.

 

UPDATE: (6/09/2017)

The Aura Tropospheric Emission Spectrometer (TES) instrument experienced another stall condition on March 12, 2017 shortly after regaining a fully operational metrology laser.  Recovery activities were ongoing from when the stall occurred.  On May 28, 2017, the TES Interferometer Control System (ICS) translator regained full motion. TES then resumed science collection on May 30, 2017 with both full travel on the translator and a working metrology laser.

 

UPDATE: (2/18/2017)

The TES ground data system was updated with a patch (Release 14.1) on February 18, 2017 to the Operational SIPS.  This enhancement will provide improved TES Level 2 H2O, CO2 and atmospheric temperature retrievals.  Additionally, the V007 Instantaneous Radiative Kernel (IRK) L2 product will now be generated and delivered to the DAAC.  Reprocessing of the previously delivered Release 14 products will be required but shouldn’t significantly impact the final delivery date for the entire Release 14 reprocessing effort.

 

UPDATE: (1/24/2017)

 The TES instrument metrology laser end of life testing that resumed on January 11, 2017 provided a gradual increase in the laser diode current. This increase has resulted in the restoration of the primary metrology laser.  The TES instrument was transitioned back to laser metrology mode on January 24 and laser A is again producing science interferograms.   The operational TES instrument data collection has been reconfigured to use the metrology laser and SimClk (on-board clock simulator) to continue acquisition of raw atmospheric species data.  The metrology data collected data is being processed through the entire ground data system and is being routinely delivered to the DAAC as part of the Release 14 software reprocessing effort.   The TES project team is continuing the SimClk software adaptation and anticipates that processed L1B and L2 data will be available in 2017.

 

UPDATE: (1/11/2017)

The TES instrument metrology laser use was discontinued (due to laser failure) in July 2016.  The operational TES instrument data collection has been reconfigured to use the SimClk (on-board clock simulator) to continue acquisition of raw atmospheric species data.  The data is currently being processed through L1A and delivered to the DAAC.  Algorithm and engineering activities are on-going to update the L1B and Level 2 processing algorithms to accommodate this new data collection methodology.  The TES project team anticipates that processed L1B and L2 data will be available in September/October 2017.

 

The TES project is continuing with the Release 14 software reprocessing effort and is delivering L1B (version 006) and L2 (version 007) data products to the ASDC.

 

 

UPDATE:  The TES ICS translator recovered within less than 24 hours after the Aug. 15, 2016 stall and was placed in "SimClock" since Tuesday (8/16/16).

UPDATE:  The TES ICS translator has again stalled.  The stall occurred during runcount 28631 at 09:43:09 GMT on Monday (8/15/16).

UPDATE:  The TES ICS translator recovered within less than 24 hours after the Aug. 4, 2016 stall and was placed in "SimClock" since Monday (8/8/16).

UPDATE:   The TES ICS translator has stalled a third time in as many weeks.  The stall occurred at 06:07 GMT on Aug 4, 2016. More information will be provided as it becomes available.

UPDATE:   A 3-day data outage involving the ICS took place from July 20, 2016 through July 23, 2016. Special Observations resumed on July 24, 2016.

 

On Wednesday, July 20, 2016 the EOS-Aura Tropospheric Emission Spectrometer (TES) instrument operations were suspended due to issues with the interferometer control system. Other TES subsystems remain operational, and TES temperatures and currents are nominal. Based on previous stall events, full recovery is expected to take a few days to weeks. Additional details on the recovery activities will be provided as they become available.